# Two-scale Neural Networks for Partial Differential Equations with Small Parameters

### v1

@@ Abstract.
Keywords: two-scale neural networks, partial differential equations, small parameters
@@ 1. Introduction
(1.1) $d\partial_{t}u-\varepsilon L_{2}u+L_{0}u=f,\quad\mathbf{x}\in D\subset\mathbb{R}^{d},$
where some proper boundary conditions and initial conditions are imposed. Here $d=0$ or $1$ and $\varepsilon>0$. Also, $L_{2}u$ consists of the leading-order differential operator and $L_{0}u$ consists of lower-order linear or nonlinear differential operators. For example, consider a singular perturbation problem where $d=0$, $L_{2}u={\rm div}(a\nabla u)$ is second-order and $L_{0}u=b\cdot\nabla u+cu$ is first-order.
Small parameters in the equation often pose extra difficulties for numerical methods, see e.g., [27]. The difficulties come from one or more sharp transitions in regions of small volumes, which implies large first-order derivatives or even large high-order derivatives. When $d=0$, $\varepsilon>0$ is very small, $L_{2}u=\Delta u$ and $L_{0}u=\mathbf{1}\cdot\nabla u$, then at least one boundary layer arise.
When deep feedforward neural networks are used, they are usually trained with stochastic gradient descent methods but do not resolve the aforementioned issues as they learn functions with low frequency and small first-order derivatives, see e.g., in [3, 26, 38].
@@ 1.1. Literature review
  • Enhanced loss by adding first-order and higher-order derivatives: Another approach is to include the gradient information in the loss function, e.g., in [6, 12, 17]. This approach has been applied to solve partial differential equations, e.g., in [28, 40].
@@
Other approaches are not particularly for the high-frequency issue, such as [14] using adaptive activation functions and [32] by respecting causality.
@@ 1.3. Contribution and significance of the work
We directly incorporate the scale parameter into the neural networks’ architecture, establishing a two-scale neural network method. In particular, for a small scale parameter $\epsilon$, the networks employ auxiliary variables related to the scale of ${\varepsilon}^{\gamma}$ ($\gamma\in\mathbb{R}$) designed to capture intricate features that involve large derivatives such as boundary layers, inner layers, and oscillations. Specifically, we use the feedforward neural network of the form
(1.2) $w(\mathbf{x})N(\mathbf{x},\varepsilon^{\gamma}(\mathbf{x}-\mathbf{x}_{c}),\varepsilon^{\gamma}),$
@@
(1.3) $w(\mathbf{x})N(\mathbf{x}).$
@@ 2. Formulation and algorithms
$\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\left|r\left(t_{r}^{i},\boldsymbol{x}_{r}^{i}\right)\right|^{2}+\frac{\alpha}{N_{b}}\sum_{i=1}^{N_{b}}\left|u\left(t_{b}^{i},\boldsymbol{x}_{b}\right)-g^{i}\right|^{2}+\frac{\beta}{N_{0}}\sum_{i=1}^{N_{0}}\left|u\left(t_{0},\boldsymbol{x}_{0}^{i}\right)-u_{0}^{i}\right|^{2}+\frac{\alpha_{1}}{N_{c}}\sum_{i=1}^{N_{c}}|\nabla u(t_{r}^{i},\boldsymbol{x}_{r}^{i})|^{2},$
@@
In this paper, we utilize the Adam [16] optimizer to train the neural network parameters. To enhance the accuracy of capturing solutions to problems with small parameters $\epsilon$, we employ a successive training strategy. This approach progressively optimizes the weights and biases of the neural networks, starting with modestly larger $\epsilon_{0}$ as the initial guess. We summarize the successive training method in Algorithm 1.
@@
Step 1. Pick a large $\epsilon_{0}$ if $\epsilon$ is small; otherwise set $\epsilon_{0}=\epsilon$. Train optimized weights and bias of the neural networks with the scale $1$ and $\epsilon_{0}^{-\frac{1}{2}}$. Step 2. Repeat Step 1 once if necessary, e.g. $10\epsilon$ is still small. Step 3. Compute the gradients of the trained neural networks in Steps 1 & 2 on a different set of sampling points, in addition to training points. Step 4. (Adaptive sampling). Find the first few points where the norm of the gradient is large. Add a few more points. Step 5. (Fine-tuning). Use Adam with a smaller learning rate to further optimize the weights and biases. Step 6. Stop the process if the iteration attains the maximal epoch number specified.
@@ 3. Numerical results
In this section, we employ the two-scale neural network method for extensive numerical tests on problems where small parameters lead to solutions with large derivatives. The numerical tests cover one-dimensional (1D) ODEs with one (Example 3.1) and two (Example 3.2) boundary layers, 1D viscous Burgers equations with an inner layer (Example 3.3), two-dimensional (2D) steady-state convection-diffusion problems featuring two boundary layers (Example 3.4), and 2D Helmholtz problems characterized by oscillations acting as inner layers (Example 3.5). The numerical results indicate that the two-scale neural network method can provide reasonable accuracy in capturing features arising from large derivatives in solutions. These features include boundary layers, inner layers, and oscillations.
For all the numerical tests, we let $\gamma=-\frac{1}{2}$ ($\gamma$ is defined in (1.2)). Without stating otherwise, we adopt the piecewise constant learning rates scheduler, the uniform distribution for collocation points, and the $\tan h$ activation function, as listed in Table 1. For some numerical tests, we employ the successive training strategy listed in Algorithm 1.
@@
Table 1. Learning rates scheduler, collocation points distribution, and activation function
@@
$-\varepsilon u^{\prime\prime}+2u^{\prime}=3,\quad\quad u(0)=0,\quad u(1)=0.$
The exact solution to this problem is
@@
The solution has a boundary layer at $x=1$. We use the two-scale neural networks $N(x,(0.5-x)/\sqrt{\epsilon},1/\sqrt{\epsilon})$ to solve the ODE problem. The neural network (NN) size is $(3,20,20,20,20,1)$. We collect the numerical results in Figures 1 and 4(a) for the case where $\epsilon=10^{-2}$. Observing from Figure 1(b) to (d), the NN solution matches the exact solution very well, with the relative errors around the boundary layer $x=1$ only at the magnitude of $10^{-3}$.
For the case where $\epsilon=10^{-3}$, we collect the results in Figures 2 and 4(b). As depicted in Figure 2(a), the NN solution captures the behavior of the exact solution. While there is a relatively larger error observed very close to $x=1$ as shown in Figure 2(b) and (c), it is essential to note that this challenge is inherent to the nature of the problem itself. In particular, when $\epsilon=10^{-3}$, the boundary layer essentially takes the form of a vertical line. This results in a heightened sensitivity to the shift between the actual and predicted boundary layer in the vicinity of $x=1$, as shown in Figure 2(c).
@@
Figure 1. numerical results for Example 3.1 when $\epsilon=10^{-2}$ using $N(x,(0.5-x)/\sqrt{\epsilon},1/\sqrt{\epsilon})$, $\alpha=1,\alpha_{1}=0$, $N_{c}=200$, epoch=20000.
@@
Figure 2. numerical results for Example 3.1 when $\epsilon=10^{-3}$, using $N(x,(0.5-x)/\sqrt{\epsilon},1/\sqrt{\epsilon})$, $\alpha=1000$, $\alpha_{1}=10^{-6}$, $N_{c}=450$, epoch=25000.
@@
(3.1) $\varepsilon u^{\prime\prime}-xu^{\prime}-u=0,\quad-1<x<1,\quad u(-1)=1,\quad u(1)=2.$
The exact solution to the problem is
@@
The solution has two boundary layers, at $x=\pm 1$.
We solve the considered problem using the two-scale NN $N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$. The NN size is $(3,20,20,20,20,1)$. We collect the numerical results in Figures 3 and 4(c) for the case when $\epsilon=10^{-2}$. The figures illustrate the excellent agreement between the NN solution and the exact solution, even at the boundary layers. As shown in Figure 3(c) and (d), the relative error is at the level of $10^{-2}$ in the vicinity of the left boundary layer and around $10^{-3}$ near the right boundary layer. These results underscore the capability of the two-scale NN method in capturing multiple boundary layers in the exact solution.
For the scenario when $\epsilon=10^{-3}$, the boundary layers practically become vertical lines, as shown in Figure 5(a). To address this inherent difficulty, we adopt the successive training strategy outlined in Algorithm 1. To be specific, we initialize the parameters (weights and biases) of $N\left(x,\frac{x}{\sqrt{\epsilon}},\frac{1}{\sqrt{\epsilon}}\right)$ when $\epsilon=10^{-3}$ using the pre-trained parameters obtained from training with $\epsilon=10^{-2}$. We collect the results in Figure 5. While the NN solution does not achieve the same level of precision as when $\epsilon=10^{-2}$, it is crucial to note that this discrepancy is primarily due to the inherent difficulty of the problem itself: the two boundary layers essentially become vertical lines. Despite the challenge, the two-scale NN solution still captures the key features of the exact solution, as shown in Figure 5(a). In addition, we provide a visualization of the shift between the actual and predicted solutions in Figure 5(c) and (d), illustrating that the shift occurs within a very narrow band around the right boundary.
@@
Figure 3. numerical results for Example 3.2 when $\epsilon=10^{-2}$, using $N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$, $\alpha=1,\alpha_{1}=0$, $N_{c}=500$, epoch=20000.
@@
Figure 4. left: full and truncated loss history for Example 3.1 when $\epsilon=10^{-2}$ and $\epsilon=10^{-3}$ , and for Example 3.2 when $\epsilon=10^{-2}$. The configurations for these three cases are described in Figures 1, 2, and 3, respectively.
@@
Figure 5. numerical results for Example 3.2 when $\epsilon=10^{-3}$ using $N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$, $\alpha=1$, $\alpha_{1}=10^{-4}$ when $\epsilon=10^{-3}$, $N_{c}=430$, epoch=50000. The parameters of the NN are pre-trained using $N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$ with $\epsilon=10^{-2}$, with constant learning rate $10^{-3}$, $\alpha=1,\alpha_{1}=0$, $N_{c}=500$, epoch=30000.
@@
(c) full and truncated loss history when $x_{0}=0$
@@
Figure 6. numerical results for Example 3.3 when ${\epsilon=10^{-2}/\pi}$ using $(x^{2}-1)N(t,x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})+g(t)$ at $t=1$. $\alpha=1,\alpha_{1}=0$, $N_{c}=22500$ on the plane $(t,x)$, epoch=60000, with constant learning rate $10^{-4}$ (pre-trained with $\epsilon=10^{-1}/\pi$, epoch=30000). The relative error is set to zero at $x=-1,0,1$ (when $x_{0}=0)$ and at $x=\pm 0.5$ (when $x_{0}=0.5$), where $|u(1,x)|<10^{-14}$.
@@
First, we discuss the case when $\epsilon=10^{-2}/\pi$. For $x_{0}=0$, i.e., the inner layer (nearly a shock profile) is located at the center of the domain. According to (3.3), $x_{0}=0$ leads to $g(t)=0$. We employ the successive training strategy in Algorithm 1 by initializing the parameters of this NN using the pre-trained parameters obtained from training with $\epsilon=10^{-1}/\pi$. We collect the numerical results in Figure 6. From Figure 6(e), we can see that the maximal relative error at the final time node $t=1$ is 2.5%. The absolute error around the inner layer, i.e., around $x=0$, is only at the level of $10^{-3}$, which accounts for a mere 0.4% of the amplitude of the inner layer.
@@
When $\epsilon=10^{-3}/\pi$, it is difficult to achieve accurate numerical quadrature for the exact solution in (3.3) because the integrand can vary from tiny scales to infinity, driven by the very small $\epsilon$. Therefore, we employ the high-resolution finite element solution as the reference solution. The NN solution is obtained through a two-phase successive training process with Algorithm 1. To be specific, we initialize the NN parameters with the pre-trained parameters obtained from training with $\epsilon=10^{-2}/\pi$. These pre-trained parameters themselves were initially obtained from the training process with $\epsilon=10^{-1}/\pi$. As observed in Figure 7, there is a shift between the exact solution and the predicted solution by the two-scale NN method within a very narrow band around the inner layer at $x=0$. The observed deviation primarily arises from the inherent difficulty of the problem itself. Indeed, given $\epsilon=10^{-3}/\pi$, the inner layer practically becomes a vertical line. Nevertheless, the two-scale NN solution still captures the key features of the exact solution in this context.
Figure 7. left: high resolution FEM solution with mesh size 1/500, polynomial degree 2 and NN solution obtained with $N(t,x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$ for Example 3.3 when $\epsilon=10^{-3}/\pi,x_{0}=0$ at $t=1$. $\alpha=1$, $N_{c}=22500$ on the plane $(t,x)$, epoch=60000 (two-phase successive training: pre-trained with $\epsilon=10^{-2}/\pi$, epoch=60000; the pre-trained parameters are obtained from training with $\epsilon=10^{-1}/\pi$, epoch=30000); right: relative error.
@@
(f) full and truncated loss history
(g) absolute error around $x=1$, $y=1$. (top view)
(h) relative error around $x=1$, $y=1$. (top view)
Figure 8. numerical results for Example 3.4 when $\epsilon=10^{-2}$ using $N(x,(0.5-x)/\sqrt{\epsilon},1/\sqrt{\epsilon},y,(0.5-y)/\sqrt{\epsilon},1/\sqrt{\epsilon})$ with $\alpha=100$, $\alpha_{1}=0,N_{c}=22500,$ epoch=35000. We enforce $e_{rel}(x,y)=0$ on $\partial\Omega$ where the exact solution $u=0$. The relative $l^{2}$ error is $e_{rel}^{l^{2}}=1.5722\times 10^{-5}.$
Example 3.5 (2D Helmoholtz problem).

### v2

@@ Abstract.
Keywords: two-scale neural networks, partial differential equations, small parameters, successive training
@@ 1. Introduction
(1.1) ${p}\partial_{t}u-\epsilon L_{2}u+L_{0}u=f,\quad\mathbf{x}\in D\subset\mathbb{R}^{d},~{}{t\in(0,T],}$
where some proper boundary conditions and initial conditions are imposed. Here ${p}=0$ or $1$, $\epsilon>0$, $D$ is the spatial domain, ${d\in\mathbb{N}^{+}}$ and $T>0$. Also, $L_{2}u$ consists of the leading-order differential operator and $L_{0}u$ consists of lower-order linear or nonlinear differential operators. For example, consider a singular perturbation problem where ${p}=0$, $L_{2}u={\rm div}({a}\nabla u)$ is second-order and $L_{0}u=b\cdot\nabla u+cu$ is first-order.
Small parameters in the equation often pose extra difficulties for numerical methods, see e.g., [29]. The difficulties come from one or more sharp transitions in regions of small volumes, which implies large first-order derivatives or even large high-order derivatives. When ${p}=0$, $\epsilon>0$ is very small, $L_{2}u=\Delta u$ and $L_{0}u=\mathbf{1}\cdot\nabla u$, then at least one boundary layer arises.
When deep feedforward neural networks are used, they are usually trained with stochastic gradient descent methods but do not resolve the issues above as they learn functions with low frequency and small first-order derivatives, see e.g., in [3, 28, 40].
@@ 1.1. Literature review
  • Enhanced loss by adding first-order and higher-order derivatives: Another approach is to include the gradient information in the loss function, e.g., in [6, 13, 19]. This approach has been applied to solve PDEs, e.g., in [30, 42].
@@
Other approaches are not particularly for the high-frequency issue, such as [16] using adaptive activation functions and [34] by respecting causality, continual learning [14] and sequential training [12].
@@ 1.3. Contribution and significance of the work
We directly incorporate the scale parameter into the neural networks’ architecture, establishing a two-scale neural network method. In particular, for a small scale parameter $\epsilon$, the networks employ auxiliary variables related to the scale of ${\epsilon}^{\gamma}$ ($\gamma\in\mathbb{R}$) designed to capture intricate features that involve large derivatives such as boundary layers, inner layers, and oscillations. Specifically, we use the feedforward neural network of the form
(1.2) $w(\mathbf{x})N({t},\mathbf{x},\epsilon^{\gamma}(\mathbf{x}-\mathbf{x}_{c}),\epsilon^{\gamma}),\quad\gamma<0,$
@@
(1.3) $w(\mathbf{x})N({t},\mathbf{x}).$
We note that when addressing steady-state problems, $t$ is removed from (1.2) and (1.3).
The network (1.2) is motivated as follows. First, it can handle large derivatives for training neural networks. By taking derivatives in $x$, we readily obtain a factor of $\epsilon^{\gamma}$. Thus, the networks (1.2) can facilitate learning large derivatives or higher frequencies, especially with low-order optimizers. Second, the network (1.2) is analogous to the classical scale expansion of $u$ in $\epsilon$ while we do not have explicit expansions. Suppose that $u=\sum_{k=0}^{\infty}u_{k}\epsilon^{k\beta}$, where $\beta>0$ is problem-dependent. The intuition is that the scale of $u_{k}$ is about the order of $\epsilon^{-k\beta}$. For the network (1.2), Taylor’s expansion in $\mathbf{x}$ leads to a scale expansion in $\epsilon$. Consider the one-dimensional function and the first two derivatives of (1.2). The first derivative is of order $\epsilon^{\gamma}$. The second derivative is of the order $\epsilon^{2\gamma}$. For singular perturbation problems in this work, $\beta$ is usually $1/2$ or $1$. We then choose $\gamma=-1/2$ in all cases. In some cases, $\gamma=-1$ may be a better choice but still can be covered if we use a smooth activation function since $(\epsilon^{-\frac{1}{2}})^{2}=\epsilon^{-1}$. It may enhance the flexibility of networks by using $\gamma=-1/2$ in all cases, especially when a smooth activation function implies a valid Taylor’s expansion in $\epsilon$.
The point $\mathbf{x}_{c}$ is chosen as the center of the spatial domain $D$. However, other choices are possible. For example, we can take $\mathbf{x}_{c}$ to be an endpoint if there is only a boundary layer at one endpoint. However, in most cases, we do not know where large derivatives are. In these cases, we expect that the linear inner layer of the feedforward neural network can help find the best linear combinations of $\epsilon^{\gamma}(\mathbf{x}-\mathbf{x}_{c})$ and $\epsilon^{\gamma}$ in training and consequently have locations of large derivatives.
@@ 2. Formulation and algorithms
(2.2) $\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\left|r\left(t_{r}^{i},\boldsymbol{x}_{r}^{i}\right)\right|^{2}+\frac{\alpha}{N_{b}}\sum_{i=1}^{N_{b}}\left|u\left(t_{b}^{i},\boldsymbol{x}_{b}\right)-g^{i}\right|^{2}+\frac{\beta}{N_{0}}\sum_{i=1}^{N_{0}}\left|u\left(t_{0},\boldsymbol{x}_{0}^{i}\right)-u_{0}^{i}\right|^{2}+\frac{\alpha_{1}}{N_{c}}\sum_{i=1}^{N_{c}}|\nabla u(t_{r}^{i},\boldsymbol{x}_{r}^{i})|^{2},$
@@
In this paper, we utilize the Adam [18] optimizer to train the neural network parameters. To enhance the accuracy of capturing solutions to problems with small parameters $\epsilon$, we employ a successive training strategy. This approach progressively optimizes the weights and biases of the neural networks, starting with modestly larger $\epsilon_{0}$ as the initial guess. We summarize the successive training method in Algorithm 1. We remark that this training strategy is similar to continual learning in [14] and sequential training in [12]. However, we fix the size of neural networks, while in [12, 14], the network size increases during different stages of training.
@@
Pick $\epsilon_{0}=O(0.1)$ if $\epsilon$ is very small. Initialize the weights and biases randomly.
Step 1. Solve the PDE with parameter $\epsilon_{0}$ instead of $\epsilon$. Use the loss function above with $\epsilon_{0}$ in place of $\epsilon$. Train to obtain weights and biases of the neural networks (1.2) with $\epsilon=\epsilon_{0}$.
Step 2. Pick a positive integer $\ell>1$. While $\epsilon_{0}\geq\ell\epsilon$
    set $\epsilon_{0}=\epsilon_{0}/\ell$ and go to Step 1;
  end
If $\epsilon_{0}<\ell\epsilon$, set $\epsilon_{0}=\epsilon$ and go to Step 1. In these iterations, initialize the weights and biases obtained from the last iteration.
Step 3. (Fine-tuning). Use Adam with a smaller learning rate or a second-order optimizer to further optimize the weights and biases.
Step 4. Stop the process if the maximal epoch number is reached.
@@ 3. Numerical results
In this section, we employ the two-scale neural network method for extensive numerical tests on problems where small parameters lead to solutions with large derivatives. The numerical tests cover one-dimensional (1D) ODEs with one (Examples 3.1, 3.6 and 3.7) and two (Example 3.2) boundary layers, 1D viscous Burgers equations with an inner layer (Example 3.3), two-dimensional (2D) steady-state convection-diffusion problems featuring two boundary layers (Example 3.4), and 2D Helmholtz problems characterized by oscillations acting as inner layers (Example 3.5). The numerical results indicate that the two-scale neural network method can provide reasonable accuracy in capturing features arising from large derivatives in solutions. These features include boundary layers, inner layers, and oscillations.
For all the numerical tests, we let $\gamma=-\frac{1}{2}$ in (1.2). Unless stated otherwise, regardless of whether Algorithm 1 is used, we adopt the default setup up of the piecewise constant learning rates scheduler, the uniform distribution for collocation points, and the ${\tanh}$ activation function, as listed in Table 1. All examples are implemented with JAX and the Adam optimizer unless stated otherwise. In some tests, we employ the successive training strategy listed in Algorithm 1. The parameters $\epsilon_{0}$, $\ell$ are manually set and are problem-dependent. We let the starting $\epsilon_{0}=0.1$ and $\ell=10$ unless stated otherwise. In Examples 3.1 and 3.2, when $\epsilon=10^{-5}$, we use $\ell=10$ for the successive training until $\epsilon_{0}\geq 10^{-4}$ and then $\ell=2$ to reach $\epsilon=10^{-5}$
To further demonstrate the robustness of numerical results, we report the statistical metric $\overline{e}\pm\sigma$ for Examples 3.1 and 3.2. Here, $\overline{e}$ is the average absolute error and $\sigma$ is the standard deviation.
@@
Table 1. Default setup (unless stated otherwise) of the learning rates scheduler, collocation points distribution, and activation function
@@
$-\epsilon u^{\prime\prime}+2u^{\prime}=3,\quad{0<x<1,}\quad u(0)=0,\quad u(1)=0.$
The solution has a boundary layer at $x=1$. The exact solution to this problem is
@@
We use the two-scale neural networks $N(x,(0.5-x)/\sqrt{\epsilon},1/\sqrt{\epsilon})$ to solve the ODE problem. The neural network (NN) size is $(3,20,20,20,20,1)$. We employ the successive training strategy in Algorithm 1 starting with $\epsilon_{0}=10^{-1}$, and successively solve the problem for $\epsilon$ as small as $10^{-5}$, following the sequence $\epsilon_{0}=10^{-1},10^{-2},10^{-3},10^{-4},5\times 10^{-5},2.5\times 10^{-5},1.25\times 10^{-5},10^{-5}$. The parameters in the discrete loss function in (2.2) and hyper-parameters of the training are summarized in Table 2 . We collect the numerical results in Figures 1 and 5(a) for the case where $\epsilon=10^{-2}$. Observing from Figure 1(b) to (d), the NN solution matches the exact solution very well, with the relative errors around the boundary layer $x=1$ only at the magnitude of $10^{-3}$.
For the case where $\epsilon=10^{-3}$, we collect the results in Figures 2 and 5(b). As depicted in Figure 2(a), the NN solution captures the behavior of the exact solution. While there is a relatively larger error observed very close to $x=1$ as shown in Figure 2(b) and (c), it is essential to note that this challenge is inherent to the nature of the problem itself. In particular, when $\epsilon=10^{-3}$, the boundary layer essentially takes the form of a vertical line. This increases sensitivity to the shift between the actual and predicted boundary layer around $x=1$, as shown in Figure 2(c). We also present the results for $\epsilon=10^{-4}$ and $10^{-5}$ in Figure 3, where we observe comparable accuracy as in the case $\epsilon=10^{-3}$.
$\epsilon$ $10^{-1}$ (starting $\epsilon_{0}$) $10^{-2}$ $10^{-3}$ $10^{-4}$ $\boldsymbol{\epsilon}_{s}$
$\alpha$ 1 100 1000 1000 1000
$\alpha_{1}$ 0 0 $10^{-6}$ $10^{-6}$ $10^{-5}$
$N_{c}$ 200 200 450 450 300
LR PC $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
epochs 20000 30000 50000 50000 30000 each
Table 2. Parameters in the loss function (2.2) and hyper-parameters of the successive training for Example 3.1. Here, $\boldsymbol{\epsilon}_{s}=10^{-5}\times[5,2.5,1.25,1]$, LR is the abbreviation for learning rate, PC is the piecewise constant scheduler in Table 1.
@@
Figure 1. Numerical results for Example 3.1 when $\epsilon=10^{-2}$ using $N(x,(x-0.5)/\sqrt{\epsilon},1/\sqrt{\epsilon})$, with parameters specified in Table 2 .
@@
Figure 2. Numerical results for Example 3.1 when $\epsilon=10^{-3}$, using $N(x,(x-0.5)/\sqrt{\epsilon},1/\sqrt{\epsilon})$, with parameters specified in Table 2 .
(a) exact and NN solutions for $\epsilon=10^{-4}$
(b) absolute error
(c) relative error around the boundary layer
(d) relative error away from the boundary layer
(e) exact and NN solutions for $\epsilon=10^{-5}$
(f) absolute error
(g) relative error around the boundary layer
(h) relative error away from the boundary layer
Figure 3. Numerical results for Example 3.1 when $\epsilon=10^{-4}$ ((a) to (d)) and $10^{-5}$ ((e) to (h)), with parameters specified in Table 2.
@@
(3.1) $\epsilon u^{\prime\prime}-xu^{\prime}-u=0,\quad-1<x<1,\quad u(-1)=1,\quad u(1)=2.$
The solution has two boundary layers, at $x=\pm 1$. The exact solution to the problem is
@@
We solve the considered problem using the two-scale NN $N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$ We employ the successive training strategy in Algorithm 1 starting with $\epsilon_{0}=10^{-1}$, and successively solve the problem for $\epsilon$ as small as $10^{-5}$, following the sequence $\epsilon_{0}=10^{-1},10^{-2},10^{-3},10^{-4},5\times 10^{-5},2.5\times 10^{-5},1.25\times 10^{-5},10^{-5}$. The parameters in the discrete loss function in (2.2) and hyper-parameters of the training are summarized in Table 3. The NN size is $(3,20,20,20,20,1)$. We collect the numerical results in Figures 4 and 5(c) for the case when $\epsilon=10^{-2}$. The figures illustrate the excellent agreement between the NN solution and the exact solution, even at the boundary layers. As shown in Figure 4 (c) and (d), the relative error is at the level of $10^{-2}$ in the vicinity of the left and right boundary layers. These results underscore the capability of the two-scale NN method in capturing multiple boundary layers in the exact solution.
[…]
$\epsilon$ $10^{-1}$ (starting $\epsilon_{0}$) $10^{-2}$ $10^{-3}$ $10^{-4}$ $\boldsymbol{\epsilon}_{s}$
$\alpha$ 1 1 1 1 1
$\alpha_{1}$ 0 0 0 $10^{-4}$ $10^{-4}$
$N_{c}$ 500 500 430 430 430
LR PC PC $10^{-4}$ $10^{-4}$ $10^{-4}$
epochs 20000 20000 50000 50000 20000 each
Table 3. Parameters in the loss function (2.2) and hyper-parameters of the successive training for Example 3.2. Here, $\boldsymbol{\epsilon}_{s}=10^{-5}\times[5,2.5,1.25,1]$, LR is the abbreviation for learning rate, PC is the piecewise constant scheduler in Table 1.
@@
Figure 4. Numerical results for Example 3.2 when $\epsilon=10^{-2}$ using $N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$, with parameters specified in Table 3.
@@
(d) $\epsilon=10^{-3}$ for Example 3.2
Figure 5. Loss and errors history, as well as the statistical metrics $\overline{e}$ and $\overline{e}\pm\sigma$ for Examples 3.1 and 3.2 when $\epsilon=10^{-2}$ and $\epsilon=10^{-3}$.
@@
Figure 6. Numerical results for Example 3.2 when $\epsilon=10^{-3}$ using $N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$, with parameters specified in Table 3.
(a) exact and NN solutions
(b) absolute error
(c) exact and NN solutions around the left boundary
(d) absolute error around the left boundary
Figure 7. Numerical results for Example 3.2 when $\epsilon=10^{-5}$ using $N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$, with parameters specified in Table 3 .
@@
(c) loss history when $x_{0}=0$
@@
Figure 8. Numerical results for Example 3.3 when ${\epsilon=10^{-2}/\pi}$ using $(x^{2}-1)N(t,x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})+g(t)$ at $t=1$. $\alpha=1,\alpha_{1}=0$, $N_{c}=22500$ on the plane $(t,x)$, epochs=60000, with constant learning rate $10^{-4}$ (pre-trained with ${\epsilon_{0}}=10^{-1}/\pi$, epochs=30000). The relative error is set to zero at $x=-1,0,1$ (when $x_{0}=0)$ and at $x=\pm 0.5$ (when $x_{0}=0.5$), where $|u(1,x)|<10^{-14}$.
@@
First, we discuss the case when $\epsilon=10^{-2}/\pi$. For $x_{0}=0$, i.e., the inner layer (nearly a shock profile) is located at the center of the domain. According to (3.3), $x_{0}=0$ leads to $g(t)=0$. We employ the successive training strategy in Algorithm 1 by initializing the parameters of this NN using the pre-trained parameters obtained from training with ${\epsilon_{0}}=10^{-1}/\pi$. We collect the numerical results in Figure 8. From Figure 8(e), we can see that the maximal relative error at the final time node $t=1$ is 2.5%. The absolute error around the inner layer, i.e., around $x=0$, is only at the level of $10^{-3}$, which accounts for a mere 0.4% of the amplitude of the inner layer.
@@
When $\epsilon=10^{-3}/\pi$, it is difficult to achieve accurate numerical quadrature for the exact solution in (3.3) because the integrand can vary from tiny scales to infinity, driven by the very small $\epsilon$. Therefore, we employ the high-resolution finite element solution as the reference solution. The NN solution is obtained through a two-phase successive training process with Algorithm 1. To be specific, we initialize the NN parameters with the pre-trained parameters obtained from training with ${\epsilon_{0}}=10^{-2}/\pi$. These pre-trained parameters themselves were initially obtained from the training process with ${\epsilon_{0}}=10^{-1}/\pi$. As observed in Figure 9, there is a shift between the exact solution and the predicted solution by the two-scale NN method within a very narrow band around the inner layer at $x=0$. The observed deviation primarily arises from the inherent difficulty of the problem itself. Indeed, given $\epsilon=10^{-3}/\pi$, the inner layer practically becomes a vertical line. Nevertheless, the two-scale NN solution still captures the key features of the exact solution in this context.
Figure 9. left: high resolution FEM solution with mesh size 1/500, polynomial degree 2 and NN solution obtained with $N(t,x,x/\sqrt{\epsilon},1/\sqrt{\epsilon})$ for Example 3.3 when $\epsilon=10^{-3}/\pi,x_{0}=0$ at $t=1$. $\alpha=1$, $N_{c}=22500$ on the plane $(t,x)$, epochs=60000 (two-phase successive training: pre-trained with ${\epsilon_{0}}=10^{-2}/\pi$, epochs=60000; the pre-trained parameters are obtained from training with ${\epsilon_{0}}=10^{-1}/\pi$, epochs=30000); right: relative error.
@@
(f) loss history
(g) errors history
(h) absolute error around $x=1$, $y=1$. (top view)
(i) relative error around $x=1$, $y=1$. (top view)
Figure 10. numerical results for Example 3.4 when $\epsilon=10^{-2}$ using $N(x,(x-0.5)/\sqrt{\epsilon},1/\sqrt{\epsilon},y,(y-0.5)/\sqrt{\epsilon},1/\sqrt{\epsilon})$ with $\alpha=100$, $\alpha_{1}=0,N_{c}=22500,$ epochs=40000. We enforce $e_{rel}(x,y)=0$ on $\partial\Omega$ where the exact solution $u=0$. The relative $l^{2}$ error is $e_{rel}^{l^{2}}=1.5722\times 10^{-5}.$
Example 3.5 (2D Helmholtz problem).